OpenStudy (anonymous):

A car moving with a speed of 108 km/h is brought to rest in a distance of 60 m. The car's acceleration is -7.5 m/s/s. How much time did the car take to stop?

OpenStudy (michele_laino):

the stopping work is given by the kinetic energy of the car, so if I call with \(F\) the stopping force, and with \(d\) the space traveled by the car, then we can write: \[Fd = \frac{1}{2}m{v^2}\] now, we have that \(F=ma\) where \(a\) is the deceleration, so we can write: \[mad = \frac{1}{2}m{v^2}\] and finally: \[d = \frac{{{v^2}}}{{2a}}\]

OpenStudy (michele_laino):

the requested time is given by the subsequent condition: \[0 = v - at\] please solve that equation for time \(t\). Of course \(v=108\;Km/h\)

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (anonymous):

Thank you!!!!!!!!